OpenStudy (anonymous):

Three fair coins are tossed. What is the possibility that two heads and one tail appear?

OpenStudy (anonymous):

Love questions like this. these are called bernoulli trials.

OpenStudy (anonymous):

3/8?

OpenStudy (anonymous):

P(1st head) = 1/2 (i.e out of 2 possible outcomes) P(2nd head) = 1/2 (i.e the first event does not effect the second event) P(tail) = 1/2 (the 3rd event is unaffected by the two previous events)

OpenStudy (anonymous):

P(1st head AND 2nd Head AND 3rd tail) = 1/2 x 1/2 x 1/2 = 1/8

OpenStudy (anonymous):

there is your answer